OpenStudy (anonymous):

2. Charles invested $1,500 in a certificate of deposit that pays a rate of 4% per year. What will be the approximate value of the CD in 5 years? [ Compound Interest Formula: A = P(1 + r/n)nt ]

OpenStudy (anonymous):

Value will be (1500)(1.04)^5 = (1500)(1.217)= $1825.50 (1.04)^5 = 1.04*1.04*1.04*1.04*1.04 = 1.217.